# Diversity of low-mass planet atmospheres in the C-H-O-N-S-Cl system with interior dissolution, non-ideality, and condensation: Application to TRAPPIST-1e and sub-Neptunes

### v2

@@ 2.2 Solubility laws
Note. — ∗Anorthite (An), Diopside (Di), Forsterite (Fo). : Range of experimental conditions used to determine the solubility laws. The oxygen fugacity, O2 is given in log10 units relative to the iron-wüstite (IW) buffer. : Combines the solubility of sulfur as sulfide (S2-) and sulfate (SO/S6+).
The atmospheric compositions of planets with molten interiors are modulated by both the volatile endowment of the planet and the solubility of these volatiles in the melt. The solubility of a gas species depends on its fugacity (or partial pressure under ideal gas conditions) in the gas phase, the composition of the melt, and sometimes additional parameters, such as the oxygen fugacity (O2) the temperature, and total pressure. Solubility laws for major atmosphere-forming species are determined from experiments that, so far, have largely been focused on Earth-like melt compositions and conducted under , conditions motivated by Earth’s mantle. These empirically determined laws often take the form of power-law or linear fits to the experimental data to relate the dissolved concentration of a species in the melt to its fugacity in the gas phase. Importantly, solubility experiments are performed for certain melt compositions under specific temperature and pressure conditions, meaning that extrapolating these laws beyond their calibrated conditions can be erroneous. The composition of the gas phase also affects the solubility by expanding or shrinking the stability of both gas and melt species in the system. For example, the dissolution of N may occur as NH, NH3, CN-, N3- and N2 moieties in silicate melts depending on the fugacities of N-, C- and H-bearing species at a given , (Dalou et al., 2019; Grewal et al., 2020), such that solubilities are not independent of atmospheric composition. Atmodeller includes a library of solubility laws from the literature for H2, H2O, CO, CO2, CH4, N2, S2, Cl2 and the noble gases He, Ne, Ar, Kr, Xe for various melt compositions and chemical systems (Table 1). We note the need for more solubility experiments of major volatiles for diverse melt compositions beyond those of the modern Earth, and under a wide range of , conditions as well as gas phase compositions.
@@ 2.3 Real gas equations of state
Atmodeller includes a library of real gas equations of state (EOS) (Table 2), which yield expressions for the volume of any given species as a function of pressure and temperature. With this information, it is possible to quantify the relationship between a species’ fugacity and its partial pressure, expressed as the fugacity coefficient , where is fugacity and is pressure:
where is the gas constant, temperature, molar volume, , and the compressibility factor . The lower bound of the integration is the standard state pressure (usually 1 bar or 1 atmosphere, depending on the thermodynamic data source), and conventionally defines an ideal state, i.e. . Typically, experimental data or numerical calculations determine or , which are then fit to an empirical or physically motivated function of temperature and pressure, such as the Compensated Redlich-Kwong EOS (CORK, Holland & Powell, 1991) or the Virial EOS (Shi & Saxena, 1992). Because the behaviour of gases under high is related to their critical temperature and pressure, often, the pressure and temperature dependence of volume is expressed relative to these critical quantities (, ); the law of corresponding states (Van der Waals, 1873). This enables regression of experimental data across similarly behaving species and thereby improves the average fit. For known pressure and temperature, fugacity coefficients and fugacities can then be computed by evaluating the integral in Equation 1. Some functional forms have convenient closed-form solutions of the fugacity, whereas others require numerical solution (see Holland & Powell, 1991, for a discussion). It is also beneficial for numerical stability to ensure the EOS are bounded and smooth, so we extrapolate the compressibility factor, , linearly according to a Virial-type expression (Appendix D).
@@
| H2 | ab-initio | Chabrier & Debras (2021) | Also mixtures with He | 0-1014 | 100- |
@@
| He | ab-initio | Chabrier & Debras (2021) | Also mixtures with H2 | 0-1014 | 100- |
@@
Note. — ∗Compensated-Redlich Kwong (CORK), Beattie-Bridgeman (B-B). : Approximate range of thermodynamic conditions, accessed by experiments or ab initio calculations, used to inform the real gas equation of state.
@@ 2.4 Condensation
The condition for condensation is met when the addition of a condensed phase; liquid or solid, to the system of gas species satisfies the extended law of mass-action equations (xLMA). We recall that by construction the xLMA accommodates both stable and unstable condensates in the same framework through the specification of stability variables, which are independent for each condensed species and solved self-consistently as part of the solution process. For implementation details the reader is referred to Leal et al. (2016) and Kitzmann et al. (2024). For numerical stability it is necessary to define a parameter that controls the minimum amount that a stable species can achieve. Following Leal et al. (2016), we adopt , where is the minimum amount of an element in the system that composes a given species. In principle, stability variables could also be applied to gas species, although since gas-phase chemistry is notably sensitive to O2—even at very low abundance—introducing a minimum cut-off for the amounts of gas species could distort results.
@@
where, by definition, for a pure phase. We do not account for the compressibilities of condensed phases (i.e., the Poynting correction is negligible). The stable condensates included in the base package are H2O(l), C(cr), -S(cr), and ClH4N(cr,l) - ammonium chloride, because these are of primary importance for assessing planetary habitability. Thermodynamic data for condensates as well as gas species are sourced from McBride et al. (2002) and Chase (1998).
@@ 3.1.1 Context and parameters
There is significant interest in probing the atmospheres of temperate Earth-sized rocky planets, such as the TRAPPIST-1 planets (Gillon et al., 2017), , to understand the nature of potentially habitable environments beyond Earth and the solar system. The earliest outgassed atmospheres form during and immediately following planet assembly, whereby one or several epochs of magma ocean formation is thought to have distributed volatile elements readily between the atmosphere, mantle, and, potentially the metallic core (e.g., Hirschmann, 2012; Lebrun et al., 2013; Sossi et al., 2020; Chen & Jacobson, 2022; Gu et al., 2024; Dasgupta et al., 2024; Huang et al., 2024). This establishes the initial budget of volatiles in the atmosphere and interior, which may subsequently participate in geological volatile cycling that further modulates their atmospheric abundances (e.g., Sleep & Zahnle, 2001; Krissansen-Totton et al., 2018; Hakim et al., 2021). Here, we apply Atmodeller to a young TRAPPIST-1e analogue as an example of a small rocky planet during its magma ocean phase following iron-nickel-alloy core formation, although the results are applicable broadly for any Earth-sized rocky planet. These results are used to inform subsequent Atmodeller calculations to determine the post-formation atmosphere and surface abundances of TRAPPIST-1e at its present-day equilibrium temperature (Section 3.1.3).
TRAPPIST-1e has a total mass of 4.13 kg, a mantle mass of 2.91 kg (assuming the same core-mantle mass proportion as Earth), and surface radius of 5861 km (0.9 R⊕) (Agol et al., 2021). To compute the properties of the high temperature atmosphere, we assume a surface temperature of 1800 K, We consider 14 gas species: This encompasses the majority of the stable, gaseous molecules of the so-called CHONS group of elements, in addition to Cl, that constitute the majority of living matter. .
Of the gas species listed above, the following solubility laws are used to distribute them between the atmosphere and the magma ocean: H2O (Dixon et al., 1995), H2 (Hirschmann et al., 2012), CO (Yoshioka et al., 2019), CO2 (Dixon et al., 1995), CH4 (Ardia et al., 2013), N2 (Libourel et al., 2003), S-bearing species (Boulliung & Wood, 2022, 2023), Cl-bearing species (Thomas & Wood, 2021). Both the EOS of pure species and the mixing between them are assumed to be ideal. Condensed carbon (graphite) is also allowed to exist depending on thermodynamic stability, since this high temperature condensate, if present, can buffer the fugacities of carbon-bearing species at the most reducing conditions.
@@
Therefore, we fix the total abundances of N, S, and Cl based on BSE estimates, scaled to the mass of TRAPPIST-1e. The average BSE estimates for N, S, and Cl are 7.01 kg, 9.47 kg, and 1.09 kg, respectively (Sakuraba et al., 2021; Hirschmann, 2016; Kendrick et al., 2017), and, when scaled to the mantle mass of TRAPPIST-1e, they are 4.85 kg, 6.55 kg, and 7.57 kg, respectively. In a Monte Carlo simulation with 10,000 realisations we uniformly sample, in log10 space, the total amount of hydrogen between 0.1 and 10 Earth oceans , C/H (by mass) between 0.1 and 10, and the oxygen fugacity by 5 log10 units either side of the iron-wüstite (IW) buffer (Appendix E). All simulation data are available to download (Bower, 2025).
@@ 3.1.2 Volatile partitioning during the magma ocean stage
We consider TRAPPIST-1e with a fully molten mantle, for which solubility and gas phase equilibria are both important in determining the atmospheric mass and composition. For comparison, Appendix B presents the same analysis assuming a partially molten mantle with 10% melt fraction. For a fully molten mantle, solubility conspires to restrict the diversity of atmospheres (Figure 1) compared to the low melt fraction case (Figure B1). The data are coloured according to the oxygen fugacity relative to the IW buffer (IW), which defines whether the atmosphere is reduced (lower values) or oxidised (higher values). The scatter plots capture the diversity of atmospheres, whereas the contoured density levels and marginal distributions reveal the most probable range of values . Figure 1 and Figure 2 show major and minor species, respectively.
Reduced atmospheres are dominated (volume mixing ratio, VMR ) by either CO at low total H budgets or by H2 at high total H budgets, because neither carbon monoxide nor molecular hydrogen are particularly soluble in silicate melts (Hirschmann et al., 2012; Yoshioka et al., 2019). Furthermore, H2 can be accompanied by modest amounts (up to 15 mol.%) of CH4, where CH4 is buffered by graphite precipitation:
@@
However, because H2O(g) is more soluble in silicate melt than H2(g) or CO2(g), it never becomes a dominant species in the atmosphere, even under oxidising conditions (see also Bower et al., 2022; Sossi et al., 2023). Consequently, the abundances of hydrogen-bearing species in the atmosphere remain low, with H2O(g) and H2(g) comprising just a few percent of the total. Therefore, owing to its relatively lower solubility, CO2(g) assumes the role of the prevailing gas species in atmospheres between IW=+1 and IW=+3. At the highest O2s above IW=+3, SO2(g) overtakes CO2(g) as the prevailing gas species for BSE-like S abundances (see also Gaillard et al., 2022; Maurice et al., 2024; Gillmann et al., 2024). Because no other gaseous species form in appreciable quantities at these conditions, the mixing ratios of CO2(g) and SO2(g) are anti-correlated in oxidising atmospheres.
@@
The dominant outcome of preferential H2O dissolution is to enhance the relative abundances of carbon-bearing species in the atmosphere (Figure 3). This is illustrated by the mean molar mass () frequency distribution (Figure 1, bottom right), which shows a low plateau between the molar masses corresponding to H2 and H2O, a dominant peak around that of CO, a smaller peak around the molar mass of CO2, and a tail extending to SO2. The small peak at 7–8 g/mol corresponds to the most reducing, CH4–H2 atmospheres. For ‘neutral’ O2 atmospheres 1–2 log units either side of the IW buffer, the near-quantitative removal of H2O from the atmosphere means that peaks in the molar mass frequency distribution can be ascribed to a singular, dominant gas species (Figure 1, bottom row). Therefore, an atmosphere with g/mol must be CO-rich, rather than a mixture of lighter hydrogen-bearing and heavier carbon-bearing species, because H2O(g) is never a major gas species and CO2–H2 atmospheres are excluded thermodynamically (see also Sossi et al., 2023). The most oxidising atmospheres have molar masses that are defined by binary mixtures of CO2 and SO2. As such, the mean molar mass of an atmosphere is broadly positively correlated with its O2, where atmospheres with g/mol must be both reducing and H-rich (Figure 4, bottom left).
@@
A fully molten mantle always dissolves at least 60% of the total hydrogen budget. At IW close to high C/H atmospheres have marginally less total H dissolved, all else being equal, due to the stability of CH4(g) over H2(g). For all but these most reduced conditions, the dissolved fraction of H approaches 100% (Figure 4). This is evident by the sharp peak around 100% in the marginal distribution of dissolved hydrogen where even the 10th percentile exceeds 97% (see also Sossi et al., 2020). As a result, atmospheric C/H ratios can be around 4 orders of magnitude larger than total C/H, with the disparity increasing to more oxidised conditions due to increasing H2O/H2. Nearly no C is dissolved in the mantle under reducing conditions (owing to the low solubilities of CO(g) and CH4(g)), but this value increases up to about 25% of the total carbon inventory at the most oxidised conditions where CO2(g) is stable (Figure 4). This proportion decreases for correspondingly lower melt fractions (Figure B3).
@@
Total pressures range between 100 to 104 bar (90% lie within 5–500 bar), with atmospheres produced in equilibrium at intermediate O2s near the IW buffer delimiting both the high- and low ends of the range. The total pressure is largely governed by the number of ocean masses of H, together with the C/H ratio, with high C/H resulting in the highest atmospheric pressures at a given O2. Even though the mantles in equilibrium with reducing atmospheres store proportionally less H and C than their more oxidising counterparts, their lower mean molecular masses () result in only modest total pressures (Equation F1). At high O2s near IW=+4, the increasing degree of CO2 dissolution into the mantle restricts the total pressure of SO2-rich atmospheres to around 200 bar. Sulfur is almost entirely dissolved below IW=+3, but above this value up to 60% of the sulfur inventory can exist in the atmosphere (Figure 4). The non-monotonic solubility of S arises due to the occurrence of two moieties in silicate liquids; S2- below IW=+4 and SO above this value (O’Neill & Mavrogenes, 2022; Boulliung & Wood, 2023). Because the S content of the silicate liquid is proportional to O (for constant temperature, S2 and FeO content) when S2- is stable, and to O when SO is stable, the solubility of S exhibits a minimum at S2-/SO = 0.5 (where is mole fraction), or roughly IW=+4 (see also Hughes et al., 2023). Therefore, SO2 should decrease in atmospheres that are more oxidising and more reducing than those formed at the solubility minimum.
@@
Minor species for a 100% molten mantle are shown in Figure 2 (see Figure B2 for the 10% molten equivalent). Of these, N2(g) can become a major species (i.e., its mixing ratio exceeds 0.5) in a handful of cases. These atmospheres correspond to those with the lowest total pressures. That is, since the abundances of N, Cl and S are held constant (while those of H, C and O vary) in the Monte Carlo simulations, N2-rich atmospheres arise when the prescribed initial abundances of H and C are low. This behaviour is also reflected in the inverse correlation between Cl2(g) mixing ratio (which never exceeds 1.510-7) and total pressure (Figure 2). Indeed, the mixing ratios of N2(g) and Cl2(g) are positively correlated. The other minor species diverge from this expectation, because their partial pressures are sensitive to the presence of other volatile species, whose fugacities depend on O2 and/or the bulk chemical composition of the atmosphere.
@@
The second most abundant minor species is S2(g), reaching mixing ratios up to 3% (0.03), with these cases corresponding exclusively to oxidising atmospheres. As detailed above, more reducing conditions promote the dissolution of S as S2- in the silicate liquid, thereby restricting the S budget of the atmosphere and thus its ability to stabilise S2(g). A third S-bearing gas species, H2S(g), is less abundant than S2(g) and SO2(g) in atmospheres more oxidised than IW=+2, yet it becomes the dominant S-bearing gas species (mixing ratios up to 0.005) at O2s below this value. This reflects the increase in H2 at the expense of H2O, promoting the formation of H2S(g) by:
@@
The higher mixing ratios of NH3(g) (up to 0.0001) and HCl(g) (up to 0.015) in reduced atmospheres relative to their oxidised counterparts (Figure 2) are explained by reactions analogous to Equation 5 occurring when substituting S2 for N2 and for Cl2, respectively. Since the abundance of HCl on a molar basis in our simulations, HCl(g) is the dominant Cl-bearing gas under all conditions. We note, however, the lack of experimentally determined solubilities for Cl, N, and S in H-bearing systems. The mixing ratios in the 10% molten cases show generally similar behaviour, though the more soluble elements (Cl and S) exhibit higher atmospheric mixing ratios compared to thòse determined in the 100% molten cases, owing to the declining mass of molten material capable of storing them in the mantle. The corollary is that the more insoluble species (e.g., N2(g)) have relatively lower mixing ratios.

### v3

@@ 2.2 Solubility laws
Note. — ∗Anorthite (An), Diopside (Di), Forsterite (Fo). : Range of experimental conditions used to determine the solubility laws. The oxygen fugacity, O2 is given in log10 units relative to the IW buffer. : Combines the solubility of sulfur as sulfide (S2-) and sulfate (SO/S6+).
The atmospheric compositions of planets with molten interiors are modulated by both the volatile endowment of the planet and the solubility of these volatiles in the melt. The solubility of a gas species depends on its fugacity (or partial pressure under ideal gas conditions) in the gas phase, the composition of the melt, and sometimes additional parameters, such as the oxygen fugacity (O2), temperature, and total pressure. Solubility laws for major atmosphere-forming species are determined from experiments that, so far, have largely been focused on Earth-like melt compositions and conducted under , conditions motivated by Earth’s mantle. These empirically determined laws often take the form of power-law or linear fits to the experimental data to relate the dissolved concentration of a species in the melt to its fugacity in the gas phase. Importantly, solubility experiments are performed for certain melt compositions under specific temperature and pressure conditions, meaning that extrapolating these laws beyond their calibrated conditions can be erroneous. The composition of the gas phase also affects the solubility by expanding or shrinking the stability of both gas and melt species in the system. For example, the dissolution of N may occur as NH, NH3, CN-, N3- and N2 moieties in silicate melts depending on the fugacities of N-, C- and H-bearing species at a given , (Dalou et al., 2019; Grewal et al., 2020), such that solubilities are not independent of atmospheric composition. Atmodeller includes a library of solubility laws from the literature for H2, H2O, CO, CO2, CH4, N2, S2, Cl2 and the noble gases He, Ne, Ar, Kr, Xe for various melt compositions and chemical systems (Table 1). We note the need for more solubility experiments of major volatiles for diverse melt compositions beyond those of the modern Earth, and under a wide range of , conditions as well as gas-phase compositions.
@@ 2.3 Real gas equations of state
Atmodeller includes a library of real gas EOS (Table 2), which yield expressions for the volume of any given species as a function of pressure and temperature. With this information, it is possible to quantify the relationship between a species’ fugacity and its partial pressure, expressed as the fugacity coefficient , where is fugacity and is pressure:
where is the gas constant, temperature, molar volume, , and the compressibility factor . The lower bound of the integration is the standard state pressure (usually 1 bar or 1 atmosphere, depending on the thermodynamic data source), and conventionally defines an ideal state, i.e. . Typically, experimental data or numerical calculations determine or , which are then fit to an empirical or physically motivated function of temperature and pressure, such as the compensated Redlich-Kwong EOS (CORK, Holland & Powell, 1991) or the virial EOS (Shi & Saxena, 1992). Because the behavior of gases under high is related to their critical temperature and pressure, often, the pressure and temperature dependence of volume is expressed relative to these critical quantities (, ), i.e., the law of corresponding states (Van der Waals, 1873). This enables regression of experimental data across similarly behaving species and thereby improves the average fit. For known pressure and temperature, fugacity coefficients and fugacities can then be computed by evaluating the integral in Equation 1. Some functional forms have convenient closed-form solutions of the fugacity, whereas others require numerical solution (see Holland & Powell, 1991, for a discussion). It is also beneficial for numerical stability to ensure the EOS are bounded and smooth, so we extrapolate the compressibility factor, , linearly according to a virial-type expression (Appendix D).
@@
| H2 | ab initio | Chabrier & Debras (2021) | Also mixtures with He | 0-1014 | 100- |
@@
| He | ab initio | Chabrier & Debras (2021) | Also mixtures with H2 | 0-1014 | 100- |
@@
Note. — ∗Compensated-Redlich Kwong (CORK), Beattie-Bridgeman (B-B). : Approximate range of thermodynamic conditions, accessed by experiments or ab initio calculations, used to inform the real gas EOS.
@@ 2.4 Condensation
The condition for condensation is met when the addition of a condensed phase, liquid or solid, to the system of gas species satisfies the xLMA equations. We recall that by construction the xLMA accommodates both stable and unstable condensates in the same framework through the specification of stability variables, which are independent for each condensed species and solved self-consistently as part of the solution process. For implementation details, the reader is referred to Leal et al. (2016) and Kitzmann et al. (2024). For numerical stability, it is necessary to define a parameter that controls the minimum amount that a stable species can achieve. Following Leal et al. (2016), we adopt , where is the minimum amount of an element in the system that composes a given species. In principle, stability variables could also be applied to gas species, although since gas-phase chemistry is notably sensitive to O2—even at very low abundance—introducing a minimum cut-off for the amounts of gas species could distort results.
@@
where, by definition, for a pure phase. We do not account for the compressibilities of condensed phases (i.e., the Poynting correction is negligible). The stable condensates included in the base package are H2O(l), C(cr), -S(cr), and ClH4N(cr,l)—ammonium chloride, because these are of primary importance for assessing planetary habitability. Thermodynamic data for condensates as well as gas species are sourced from McBride et al. (2002) and Chase (1998).
@@ 3.1.1 Context and parameters
There is significant interest in probing the atmospheres of temperate Earth-sized rocky planets, such as the TRAPPIST-1 planets (Gillon et al., 2017), , to understand the nature of potentially habitable environments beyond Earth and the solar system. The earliest outgassed atmospheres form during and immediately following planet assembly, whereby one or several epochs of magma ocean formation is thought to have distributed volatile elements readily between the atmosphere, mantle, and, potentially the metallic core (e.g., Hirschmann, 2012; Lebrun et al., 2013; Sossi et al., 2020; Chen & Jacobson, 2022; Gu et al., 2024; Dasgupta et al., 2024; Huang et al., 2024). This establishes the initial budget of volatiles in the atmosphere and interior, which may subsequently participate in geological volatile cycling that further modulates their atmospheric abundances (e.g., Sleep & Zahnle, 2001; Krissansen-Totton et al., 2018; Hakim et al., 2021). Here, we apply Atmodeller to a young TRAPPIST-1e analog as an example of a small rocky planet during its magma ocean phase following iron-nickel-alloy core formation, although the results are applicable broadly for any Earth-sized rocky planet. These results are used to inform subsequent Atmodeller calculations to determine the postformation atmosphere and surface abundances of TRAPPIST-1e at its present-day equilibrium temperature (Section 3.1.3).
TRAPPIST-1e has a total mass of 4.13 kg, a mantle mass of 2.91 kg (assuming the same core-mantle mass proportion as Earth), and surface radius of 5861 km (0.9 R⊕) (Agol et al., 2021). To compute the properties of the high-temperature atmosphere, we assume a surface temperature of 1800 K, We consider 14 gas species: This encompasses the majority of the stable, gaseous molecules of the so-called CHONS group of elements, in addition to Cl, that constitute the majority of living matter. .
Of the gas species listed above, the following solubility laws are used to distribute them between the atmosphere and the magma ocean: H2O (Dixon et al., 1995), H2 (Hirschmann et al., 2012), CO (Yoshioka et al., 2019), CO2 (Dixon et al., 1995), CH4 (Ardia et al., 2013), N2 (Libourel et al., 2003), S-bearing species (Boulliung & Wood, 2022, 2023), Cl-bearing species (Thomas & Wood, 2021). Both the EOS of pure species and the mixing between them are assumed to be ideal. Condensed carbon (graphite) is also allowed to exist depending on thermodynamic stability, since this high-temperature condensate, if present, can buffer the fugacities of carbon-bearing species at the most reducing conditions.
@@
Therefore, we fix the total abundances of N, S, and Cl based on BSE estimates, scaled to the mass of TRAPPIST-1e. The average BSE estimates for N, S, and Cl are 7.01 kg, 9.47 kg, and 1.09 kg, respectively (Sakuraba et al., 2021; Hirschmann, 2016; Kendrick et al., 2017), and, when scaled to the mantle mass of TRAPPIST-1e, they are 4.85 kg, 6.55 kg, and 7.57 kg, respectively. In a Monte Carlo simulation with 10,000 realizations we uniformly sample, in log10 space, the total amount of hydrogen between 0.1 and 10 Earth oceans , C/H (by mass) between 0.1 and 10, and the oxygen fugacity by 5 log10 units either side of the IW buffer (Appendix E). All simulation data are available to download (Bower, 2025).
@@ 3.1.2 Volatile partitioning during the magma ocean stage
We consider TRAPPIST-1e with a fully molten mantle, for which solubility and gas-phase equilibria are both important in determining the atmospheric mass and composition. For comparison, Appendix B presents the same analysis assuming a partially molten mantle with 10% melt fraction, with major and minor species shown in Figure B1 and Figure B2, respectively. For a fully molten mantle, solubility conspires to restrict the diversity of atmospheres (Figure 1) compared to the low melt fraction case (Figure B1). The data are colored according to the oxygen fugacity relative to the IW buffer (IW), which defines whether the atmosphere is reduced (lower values) or oxidized (higher values). The scatter plots capture the diversity of atmospheres, whereas the contoured density levels and marginal distributions reveal the most probable range of values . Figure 1 and Figure 2 show major and minor species, respectively.
Reduced atmospheres are dominated (volume mixing ratio, VMR ) by either CO at low total H budgets or by H2 at high total H budgets, because neither carbon monoxide nor molecular hydrogen are particularly soluble in silicate melts (Hirschmann et al., 2012; Yoshioka et al., 2019). Furthermore, H2 can be accompanied by modest amounts (up to 15 mol%) of CH4, where CH4 is buffered by graphite precipitation:
@@
However, because H2O(g) is more soluble in silicate melt than H2(g) or CO2(g), it never becomes a dominant species in the atmosphere, even under oxidizing conditions (see also Bower et al., 2022; Sossi et al., 2023). Consequently, the abundances of hydrogen-bearing species in the atmosphere remain low, with H2O(g) and H2(g) comprising just a few percent of the total. Therefore, owing to its relatively lower solubility, CO2(g) assumes the role of the prevailing gas species in atmospheres between IW = +1 and IW = +3. At the highest O2s above IW = +3, SO2(g) overtakes CO2(g) as the prevailing gas species for BSE-like S abundances (see also Gaillard et al., 2022; Maurice et al., 2024; Gillmann et al., 2024). Because no other gaseous species form in appreciable quantities at these conditions, the mixing ratios of CO2(g) and SO2(g) are anticorrelated in oxidizing atmospheres.
@@
The dominant outcome of preferential H2O dissolution is to enhance the relative abundances of carbon-bearing species in the atmosphere (Figure 3). This is illustrated by the mean molar mass () frequency distribution (Figure 1, bottom right), which shows a low plateau between the molar masses corresponding to H2 and H2O, a dominant peak around that of CO, a smaller peak around the molar mass of CO2, and a tail extending to SO2. The small peak at 7–8 g mol-1 corresponds to the most reducing, CH4–H2 atmospheres. For ”neutral” O2 atmospheres 1–2 log units either side of the IW buffer, the near-quantitative removal of H2O from the atmosphere means that peaks in the molar mass frequency distribution can be ascribed to a singular, dominant gas species (Figure 1, bottom row). Therefore, an atmosphere with g mol-1 must be CO-rich, rather than a mixture of lighter hydrogen-bearing and heavier carbon-bearing species, because H2O(g) is never a major gas species and CO2–H2 atmospheres are excluded thermodynamically (see also Sossi et al., 2023). The most oxidizing atmospheres have molar masses that are defined by binary mixtures of CO2 and SO2. As such, the mean molar mass of an atmosphere is broadly positively correlated with its O2, where atmospheres with g mol-1 must be both reducing and H-rich (Figure 4, bottom left).
@@
A fully molten mantle always dissolves at least 60% of the total hydrogen budget. At IW close to , high C/H atmospheres have marginally less total H dissolved, all else being equal, due to the stability of CH4(g) over H2(g). For all but these most reduced conditions, the dissolved fraction of H approaches 100% (Figure 4). This is evident by the sharp peak around 100% in the marginal distribution of dissolved hydrogen where even the 10th percentile exceeds 97% (see also Sossi et al., 2020). As a result, atmospheric C/H ratios can be around 4 orders of magnitude larger than total C/H, with the disparity increasing to more oxidized conditions due to increasing H2O/H2. Nearly no C is dissolved in the mantle under reducing conditions (owing to the low solubilities of CO(g) and CH4(g)), but this value increases up to about 25% of the total carbon inventory at the most oxidized conditions where CO2(g) is stable (Figure 4). This proportion decreases for correspondingly lower melt fractions (Figure B3).
@@
Total pressures range between 100 and 104 bar (90% lie within 5–500 bar), with atmospheres produced in equilibrium at intermediate O2s near the IW buffer delimiting both the high- and low ends of the range. The total pressure is largely governed by the number of ocean masses of H, together with the C/H ratio, with high C/H resulting in the highest atmospheric pressures at a given O2. Even though the mantles in equilibrium with reducing atmospheres store proportionally less H and C than their more oxidizing counterparts, their lower mean molecular masses () result in only modest total pressures (Equation F1). At high O2s near IW = +4, the increasing degree of CO2 dissolution into the mantle restricts the total pressure of SO2-rich atmospheres to around 200 bar. Sulfur is almost entirely dissolved below IW = +3, but above this value up to 60% of the sulfur inventory can exist in the atmosphere (Figure 4). The nonmonotonic solubility of S arises due to the occurrence of two moieties in silicate liquids; S2- below IW = +4 and SO above this value (O’Neill & Mavrogenes, 2022; Boulliung & Wood, 2023). Because the S content of the silicate liquid is proportional to O (for constant temperature, S2 and FeO content) when S2- is stable, and to O when SO is stable, the solubility of S exhibits a minimum at S2-/SO = 0.5 (where is mole fraction), or roughly IW = +4 (see also Hughes et al., 2023). Therefore, SO2 should decrease in atmospheres that are more oxidizing and more reducing than those formed at the solubility minimum.
@@
Minor species for a 100% molten mantle are shown in Figure 2 (see Figure B2 for the 10% molten equivalent). Of these, N2(g) can become a major species (i.e., its mixing ratio exceeds 0.5) in a handful of cases. These atmospheres correspond to those with the lowest total pressures. That is, since the abundances of N, Cl, and S are held constant (while those of H, C, and O vary) in the Monte Carlo simulations, N2-rich atmospheres arise when the prescribed initial abundances of H and C are low. This behavior is also reflected in the inverse correlation between Cl2(g) mixing ratio (which never exceeds 1.510-7) and total pressure (Figure 2). Indeed, the mixing ratios of N2(g) and Cl2(g) are positively correlated. The other minor species diverge from this expectation because their partial pressures are sensitive to the presence of other volatile species, whose fugacities depend on O2 and/or the bulk chemical composition of the atmosphere.
@@
The second most abundant minor species is S2(g), reaching mixing ratios up to 3% (0.03), with these cases corresponding exclusively to oxidizing atmospheres. As detailed above, more reducing conditions promote the dissolution of S as S2- in the silicate liquid, thereby restricting the S budget of the atmosphere and thus its ability to stabilize S2(g). A third S-bearing gas species, H2S(g), is less abundant than S2(g) and SO2(g) in atmospheres more oxidized than IW=+2, yet it becomes the dominant S-bearing gas species (mixing ratios up to 0.005) at O2s below this value. This reflects the increase in H2 at the expense of H2O, promoting the formation of H2S(g) by:
@@
The higher mixing ratios of NH3(g) (up to 0.0001) and HCl(g) (up to 0.015) in reduced atmospheres relative to their oxidized counterparts (Figure 2) are explained by reactions analogous to Equation 5 occurring when substituting S2 for N2 and for Cl2, respectively. Since the abundance of HCl on a molar basis in our simulations, HCl(g) is the dominant Cl-bearing gas under all conditions. We note, however, the lack of experimentally determined solubilities for Cl, N, and S in H-bearing systems. The mixing ratios in the 10% molten cases show generally similar behavior, though the more soluble elements (Cl and S) exhibit higher atmospheric mixing ratios compared to those determined in the 100% molten cases, owing to the declining mass of molten material capable of storing them in the mantle. The corollary is that the more insoluble species (e.g., N2(g)) have relatively lower mixing ratios.